5(3a-1)²+3=-16(3a-1)
Answer by scott8148(6628)

let x = (3a-1)

5x^2 + 16x + 3 = 0

factoring ___ (5x + 1)(x + 3) = 0

x = -1/5 ___ 3a - 1 = -1/5 ___ a = 4/15

x = -3 ___ 3a - 1 = -3 ___ a = -2/3